Question 1.3 [4]

A set of eight cards contains one joker. Aviwe and Siphumlile are two players that are learning probability theory by playing card games with these 8 cards. Aviwe chooses 5 cards at random, Siphumlile takes the remaining three cards. What is the probability that Aviwe has the joker? Aviwe discards 4 cards and Siphumlile discards two cards. It is know that the joker has not been discarded, what is the probability that Aviwe has the joker? Explain your answer in detail.




1
Expert's answer
2021-07-26T15:25:34-0400

There are 7 cards and 1 joker. Hence a total of 8 cards.

First, Aviwe chose 5 cards at random .

The total combinations that Aviwe chooses of the 5 cards is "8C_5=56"

The total combinations that Siphumlile chooses of the 3 cards is "8C_3=56"


Probability that Aviwe has a joker is calculated as:

"\:\frac{7C_4\times 1C_1}{8C_5}\times \frac{7C_3\times 1C_0}{8C_3}=\frac{35\times 1}{56}\times \frac{35\times 1}{56}=\frac{25}{64}"

Aviwe discards 4 cards and Siphumlile discards 2 cards. Jockler has not been discarded.

Now Aviwe has 1 card and Siphumlile has 1 card.


"P\left(Aviwe\:has\:a\:Joker\right)=\frac{4C_0\times 1C_1}{5C_1}\times \frac{3C_1}{3C_1}=\frac{1}{5}\times 1=\frac{1}{5}"


Therefore, the probability that Aviwe has a Joker is "\frac{1}{5}" .